OpenStudy (anonymous):

Struggling to understand how the fundamental theorem of arithmetic should be used to prove irrationality.

OpenStudy (anonymous):

I can prove it for root 2 however something like the cube root of 7 still eludes me.

OpenStudy (anonymous):

i write cube root 7 = a/b 7=a^3/b^3 b^3*7=a^3 then im slightly stuck. there should be a contradiction but i cannot quite see it

OpenStudy (anonymous):

What is the fundamental theorem?

OpenStudy (anonymous):

every integer >1 can be expressed as a unique product of primes

OpenStudy (anonymous):

Why are you exploring this with cubes and roots? A prime is evenly divisible by only itself and 1. 1 prime 2 prime 3 prime 4 = 2*2 5 prime 6 = 3*2 Perhaps you can prove this by induction. Or by contradiction: Could two different products of primes be the same number?

OpenStudy (anonymous):

I an trying to prove that the 3rd root of 7 is irrational by contradiction.

OpenStudy (amoodarya):

\[\sqrt[3]{7}\]?

OpenStudy (amoodarya):

suppose it is rational \[\sqrt[3]{7}=\frac{ a }{ b }\rightarrow \\to \ the \ power \ of \ three\\7=\frac{ a^3 }{ b^3 }\] \[a^3=7b^3 \rightarrow a=7q\\\] now a=7q insert in formula \[(7q)^3=7b^3\\b^3=49q^3=7(7q^3)\\so\\b=7k\] b must be 7k so it is contradiction because a rational number does not simplify Q={a/b: a,b elongs to Z ,b not zero , (a,b)=1 } (a,b)=1 means a/b does not simplify like this 2/4 is not rational but 1/2 is rational

OpenStudy (anonymous):

sorry. I don't see how that used the fundamental theorem of arithmetic (FTA). I suspect a question will come up on my exam about using the FTA to prove irrationality of something but I'm never completly sure about how to do that.